# tidychangepoint: a unified framework for analyzing changepoint detection in univariate time series

### v1

@@ Abstract
We present tidychangepoint, a new R package for changepoint detection analysis. tidychangepoint leverages existing packages like changepoint, GA, tsibble, and broom to provide tidyverse-compliant tools for segmenting univariate time series using various changepoint detection algorithms. In addition, tidychangepoint also provides model-fitting procedures for commonly-used parametric models, tools for computing various penalized objective functions, and graphical diagnostic displays. tidychangepoint wraps both deterministic algorithms like PELT, and also flexible, randomized, genetic algorithms that can be used with any compliant model-fitting function and any penalized objective function. By bringing all of these disparate tools together in a cohesive fashion, tidychangepoint facilitates comparative analysis of changepoint detection algorithms and models.
@@ 1 Introduction
Let $y=\{y_{1},\ldots,y_{n}\}$ be a series of observations made over time points $t=1,\ldots,n$. Define $\tau=\{\tau_{1},\ldots,\tau_{m}\}\subseteq\{y_{t}\}$ to be a subset of the original observations known as changepoints, for some $0\leq m\leq n$. The $m$ changepoints divide the time series into $m+1$ regions, with the idea that the behavior of the time series within each region does not change, and the changepoints represent points in time at which the behavior of the underlying system changes in some meaningful way. If we adopt the convention that $\tau_{0}=1$ and $\tau_{m+1}=n+1$ then the union of the set of half-open intervals $[\tau_{j},\tau_{j+1})$ for $0\leq j\leq m$ contains the time series $y$.
For example, Figure 1 shows the difference in home runs per plate appearance between the American and National Leagues of Major League Baseball (MLB) from 1925–2022. The introduction of the designated hitter rule in the American League in 1973 led to noticeable differences in the rates at which batted balls became home runs relative to the National League (which did not adapt the rule until 2022), because the designated hitter employed by American League teams was always a better hitter than the pitcher (who was forced to bat in the National League). The years 1973 and 2022 provide known changepoints, because the data generating process changed before and after those years, with the American League expected to have higher home run rates during the intervening period. In Section 2 we will investigate how well several algorithms detect these distinct time periods.
Figure 1: Difference in the rate of home runs per plate appearance between the American and National Leagues in Major League Baseball, 1925–2022. The designated hitter rule was adopted in 1973 by the American League, but not until 2022 by the National League. Note how, during the period from 1973 to 2022 when only the American League employed a designated hitter, the difference in home run rate was nearly always negative.
@@
$\tau_{f,M,\theta_{\tau}}^{*}=\min_{\tau\subseteq\{y_{t}\}}f(\tau,M(y|\theta_{\tau}))$
@@
Thus, as applied to the baseball data mentioned above, $\tau_{BIC,meanshift,\hat{\theta}}^{*}$ is the changepoint set that minimizes the BIC as applied to the normal meanshift model with estimated parameters $\hat{\theta}$ over all possible changepoint sets. In this case, we happen to know that the true changepoint set is $\tau^{*}=\{49,98\}$, since 1973 and 2022 are the 49th and 98th entries in the time series, respectively. The true region means are $\mu=\{0.000443,-0.002475,NA\}$, where the 3rd region has no data and thus no mean. [^1] This edge case is caused by the true changepoint of 98 (i.e., 2022) being the endpoint of the time series. This is not common or interesting. The BIC value for the normal meanshift model applied to $\tau^{*}$ is $f=-875.2274$.
@@ 1.1 Existing changepoint detection algorithms
Suárez-Sierra, Coen, and Taimal (2023) detail the implementation of a changepoint detection heuristic we rebrand as Coen’s algorithm in this paper. Coen’s algorithm (Section 4.3) is genetic and uses a non-homogeneous Poisson process model (Section 5.3) to model the exceedances of a threshold over time, and a Bayesian Minimum Descriptive Length (BMDL, Section 6.3) penalized objective function. Taimal, Suárez-Sierra, and Rivera (2023) discuss modifications and performance characteristics of Coen’s algorithm.
@@ 1.2 Existing changepoint detection R packages
Likely the most prominent R package for changepoint detection is changepoint (Killick and Haynes 2022; Killick and Eckley 2014), which implements the algorithms PELT, Binary Segmentation, and Segmented Neighborhood. PELT employs one of three different models: meanshift, meanvar, and varshift (see Section 5). For the meanvar model (the most versatile) the data-generating model can follow a Normal, Poisson, Gamma, or Exponential distribution. Penalty functions (see Section 6) implemented by PELT include AIC, BIC, and MBIC, but not MDL.
The wbs (Baranowski and Fryzlewicz 2019) package implements the Wild Binary Segmentation and standard Binary Segmentation algorithms, the first of which is incorporated into tidychangepoint. The ggchangepoint (Yu 2022) package provides some wrapper functions and graphical tools for changepoint detection analysis that are similar to ours, but more limited in scope and usability. Nonparametric methods for changepoint detection are present in the ecp (N. A. James, Zhang, and Matteson 2023) and changepoint.np (Haynes and Killick 2022) packages. We consider these beyond our scope at this time, which is focused on parameteric models for changepoint detection. A Bayesian approach is implemented in the bcp (X. Wang, Erdman, and Emerson 2018; Erdman and Emerson 2008) package, which is no longer available on CRAN. The qcc (Scrucca 2017) package is more broadly focused on quality control and statistical process control, neither of which are specific to changepoint detection.
@@ 1.3 Our contribution
In this paper, we present several substantive improvements to the existing state-of-the-art. First, the tidychangepoint package provides functionality that makes working with existing changepoint detection algorithms in a tidyverse-compliant syntax easy (Section 3). This reduces friction for users already comfortable working within the tidyverse. Second, the architecture of the package is easily extended to incorporate new changepoint detection algorithms (Section 4), new parametric models (Section 5), and new penalized objective functions (Section 6). Indeed, our careful separation of these three essential elements of changepoint detection is in and of itself clarifying. To the best of our knowledge, tidychangepoint is the first R package to make genetic algorithms for changepoint detection first-class citizens, and the flexible design allows models and penalty functions to be mixed-and-matched freely. We illustrate the utility of the package in Section 2 and Section 7 and conclude in Section 8.
2 Example of tidychangepoint applied to MLB
In this section, we illustrate how tidychangepoint assists in changepoint detection analysis. As noted in Section 1, the true, known changepoints for the MLB example are 1973 and 2022.
First, we apply the PELT algorithm for a change in means and variances. This algorithm finds two changepoints, but they are not the true changepoints.
mlb_pelt <- mlb_hrs |>
  segment(method = "pelt")
changepoints(mlb_pelt, use_labels = TRUE)
[1] "$1$959-01-01"
We can try PELT again, this time specifying a model that allows only the mean to vary across regions. Unfortunately, it detects no changepoints.
mlb_pelt_meanshift <- mlb_hrs |>
  segment(method = "pelt"model_fn = fit_meanshift_norm)
changepoints(mlb_pelt_meanshift)
integer(0)
Next, we try a genetic algorithm using a normal meanshift model with autoregressed errors, as employed by Shi et al. (2022) in their analysis of temperatures in Central England. This algorithm will stop after 500 generations (maxiter), or after 50 consecutive generations with no improvement (run), whichever comes first.
mlb_ga_shi <- mlb_hrs |>
  segment(method = "ga-shi"maxiter = 500run = 50)
changepoints(mlb_ga_shi, use_labels = TRUE|>
  as_year()
[1] "1931" "1947" "1960"
This algorithm finds 3 changepoints, and creates regions of time that conform somewhat to eras traditionally demarcated by baseball analysts and historians. For example, within the time period we consider, an analysis of the metric On-Base Plus Slugging Percentage (OPS) by Woltring, Rost, and Jubenville (2018) suggests the changepoints 1941, 1960, 1976, 1993, and 2005. B. James (2014) posits the changepoints 1946, 1968, and 1992.
Finally, we employ Coen’s algorithm (Suárez-Sierra, Coen, and Taimal 2023). This algorithm also identifies no changepoints.
mlb_coen <- mlb_hrs |>
  segment(method = "ga-coen"maxiter = 500run = 50)
changepoints(mlb_coen, use_labels = TRUE)
Date of length 0
@@ 2 Example of tidychangepoint applied to MLB
Because each of the resulting objects are of class tidycpt (see Section 3), comparing the results across the four algorithms is easy.
list(mlb_pelt, mlb_pelt_meanshift, mlb_ga_shi, mlb_coen) |>
@@
1 changep~ 2.2.4   PELT      <list [1]> meanvar    MBIC       -920.   0.077 secs
2 changep~ 2.2.4   PELT      <list [1]> meanshift~ MBIC       -852.   0.203 secs
3 GA       3.2.4   Genetic   <list [1]> meanshift~ BIC        -932. 120.480 secs
4 GA       3.2.4   Genetic   <list [1]> nhpp       BMDL        153.  63.060 secs
In Figure 2, we compare the changepoint sets found by the PELT algorithm with the meanvar model and Shi’s genetic algorithm. Both algorithms created a region from 1946–1959 (or 1947–1960), and largely missed the true changepoint of 1973. Note that because the errors in Shi’s algorithm are autoregressed, the fitted values (shown in red) are not constant in each region.
@@
(b) Shi’s algorithm using the meanshift model and the BIC penalty.
Figure 2: Comparison of changepoint sets returned by two algorithms.
The remainder of this paper provides greater detail about how one can use tidychangepoint.

### v2

@@ Abstract
We present tidychangepoint, a new R package for changepoint detection analysis. Most R packages for segmenting univariate time series focus on providing one or two algorithms for changepoint detection that work with a small set of models and penalized objective functions, and all of them return a custom, nonstandard object type. This makes comparing results across various algorithms, models, and penalized objective functions unnecessarily difficult. tidychangepoint solves this problem by wrapping functions from a variety of existing packages and storing the results in a common S3 class called tidycpt. The package then provides functionality for easily extracting comparable numeric or graphical information from a tidycpt object, all in a tidyverse-compliant framework. tidychangepoint is versatile: it supports both deterministic algorithms like PELT (from changepoint), and also flexible, randomized, genetic algorithms (via GA) that—via new functionality built into tidychangepoint—can be used with any compliant model-fitting function and any penalized objective function. By bringing all of these disparate tools together in a cohesive fashion, tidychangepoint facilitates comparative analysis of changepoint detection algorithms and models.
@@ 1 Introduction
1.1 Motivating example
While global temperatures continue to rise, climate scientists continue to face baffling skepticism from some. How do we know that the increase we observe are not just anecdotal? How do we know that the recent pattern of rising temperatures is not simply the latest round of normal fluctuation?
One approach would be to measure temperature over a long period of time, and investigate the time series for the presence of changepoints. Changepoints indicate breaks in a time series that divide it into mutually exclusive regions in which the behavior of the underlying system is the same within each region, but varies across regions. A changepoint might represent a moment at which the system changed in some meaningful and measurable way.
For example, the CET data set collected by Parker, Legg, and Folland (1992) and repackaged in tidychangepoint (Baumer et al. 2024) lists the mean annual temperature in degrees Celsius from 1659 to 2024, as measured in Central England and shown in Figure 1. When did the temperature start to rise? How can we distinguish natural variation in temperature (around a static mean) from a meaningful change in that mean (caused by humans)?
library(tidychangepoint)
plot(CET)
Figure 1: Mean annual temperature in degrees Celsius, as measured in Hadley Centre, England, 1659-2024. Plot created by plot.xts().
Changepoint detection analysis is the study of such changepoints and the algorithms and models that determine them. changepoint (Killick 2024) is a popular R package that implements the PELT algorithm (Killick, Fearnhead, and Eckley 2012) for changepoint detection. We can apply this algorithm to our temperature data using the cpt.meanvar() function.
cet_cpt <- changepoint::cpt.meanvar(as.ts(CET))
In this case, PELT finds 6 changepoints, listed below and illustrated in Figure 2.
index(CET)[cet_cpt@cpts] |>
  format("%Y")
[1] "1713" "1715" "1925" "2002" "2005" "2024"
While PELT has many virtues, there are many alternative changepoint detection algorithms. Another algorithm is Wild Binary Segmentation, provided by the wbs package (Baranowski and Fryzlewicz 2024).
cet_wbs <- wbs::wbs(CET)
index(CET)[sort(cet_wbs\$cpt\$cpt.ic\$mbic.penalty)] |>
  format("%Y")
[1] "1701" "1988"
This algorithms finds 2 changepoints, as seen in Figure 2.
(a) PELT. Plot created by plot.cpt().
(b) WBS. Plot created by plot.wbs().
Figure 2: Changepoints found in temperatue data by two algorithms.
Another promising approach to determining changepoints is to use genetic algorithms that iterate over many possible changepoint sets. Shi et al. (2022) explored the use of a variety of different parametric models using genetic algorithms and a previous version of the CET data set. This analysis leveraged the GA package, which is not specific to changepoint detection, and thus each model that Shi et al. (2022) wanted to test required writing a different set of code, the complexity of which prevents us from reproducing it here. Using a trendshift model (see Section 5), they found the optimal changepoints to be 1700, 1739, and 1988.
There is no optimal solution to the Changepoint Detection Problem (CDP, defined formally in Section 1.2), and in this case, the three algorithms returned different changepoint sets, even when applied to the same data set. Furthermore—and more germaine for us—the objects returned by changepoint detection functions from changepoint, wbs, and GA are all of different types, have difffent structures, contain different information, and can be used with a variety of different functions. This heterogeneity in user interface and package architecture is widespread across roughly a dozen other changepoint detection packages we found. No standardization exists.
These incompatible software interfaces make it unnecessarily difficult to compare the performance of changepoint detection algorithms across different packages. One essentially has to write custom code for each implementation, even when using the same data set. Furthermore, since there is no pre-built interface for genetic changepoint detection algorithms, every tweak to a genetic algorithm for changepoint detection requires custom code.
While we do not claim to offer solutions to the CDP itself, tidychangepoint solves the software heterogeneity problem by providing a unified interface with built-in support for genetic algorithms for changepoint detection. Comparing algorithms and/or models for changepoint detection in tidychangepoint can be as easy as iterating over a list of specifications.
cet_objs <- list(
  null = segment(CET, method = "null"),
  pelt = segment(CET, method = "pelt"),
  wbs = segment(CET, method = "wbs"),
  shi = segment(CET, method = "ga-shi"maxiter = 1000run = 100)
Once the algorithms have been run, we can easily extract the sets of changepoints for comparison using the changepoints() function or compare the results visually using plot() (see Figure 3). Each panel of Figure 3 shows the original time series (black line) segmented by the changepoints determined by the PELT algorithm (vertical dotted lines). The horizontal red lines indicate the mean and 1.96 times the standard deviation within each region, as per the summary produced by tidy() (see Section 3.2).
cet_objs |>
  lapply(plot, use_time_index = TRUEylab = "Degrees Celsius"|>
  patchwork::wrap_plots()
Figure 3: Comparison of four different attempts to detect changepoints in the Central England temperature data. Plots created by plot.tidycpt().
cet_objs |>
  lapply(changepoints, use_labels = TRUE|>
  lapply(as_year)
\$null
character(0)
\$pelt
[1] "1713" "1715" "1925" "2002" "2005"
\$wbs
[1] "1701" "1739" "1740" "1892" "1988"
\$shi
[1] "1691" "1699" "1740" "1741" "1893" "1989"
We return to this example in Section 2.1.
1.2 Mathematical problem formulation
@@
Let $y=\{y_{1},\ldots,y_{n}\}$ be a series of observations made over time points $t=1,\ldots,n$. Define $\tau=\{\tau_{1},\ldots,\tau_{m}\}\subseteq y$ to be a subset of the original observations known as changepoints, for some $0\leq m\leq n$. The $m$ changepoints divide the time series into $m+1$ regions, with the idea that the behavior of the time series within each region does not change, and the changepoints represent points in time at which the behavior of the underlying system changes in some meaningful way. If we adopt the convention that $\tau_{0}=1$ and $\tau_{m+1}=n+1$ then the union of the set of half-open intervals $[\tau_{j},\tau_{j+1})$ for $0\leq j\leq m$ contains the time series $y$.
@@
$\tau_{f,M,\theta_{\tau}}^{*}=\min_{\tau\subseteq y}f(\tau,M(y|\theta_{\tau}))$
@@ 1.3 Existing changepoint detection algorithms
Suárez-Sierra, Coen, and Taimal (2023) detail the implementation of a changepoint detection heuristic we rebrand as Coen’s algorithm in this paper. Coen’s algorithm (Section 4.3) is genetic and uses a non-homogeneous Poisson process model (Section 5.3) to model the exceedances of a threshold over time, and a Bayesian Minimum Descriptive Length (BMDL, Section 6.1.4) penalized objective function. Taimal, Suárez-Sierra, and Rivera (2023) discuss modifications and performance characteristics of Coen’s algorithm.
@@ 1.4 Existing changepoint detection R packages
Likely the most prominent R package for changepoint detection is changepoint (Killick 2024; Killick and Eckley 2014), which implements the algorithms PELT, Binary Segmentation, and Segmented Neighborhood. PELT employs one of three different models: meanshift, meanvar, and varshift (see Section 5). For the meanvar model (the most versatile) the data-generating model can follow a Normal, Poisson, Gamma, or Exponential distribution. Penalty functions (see Section 6) implemented by PELT include AIC, BIC, and MBIC, but not MDL. EnvCpt [R-EnvCpt] is another package that wraps changepoint functionality.
The wbs (Baranowski and Fryzlewicz 2024) package implements the Wild Binary Segmentation and standard Binary Segmentation algorithms, the first of which is incorporated into tidychangepoint. The ggchangepoint (Yu 2022) package provides some wrapper functions and graphical tools for changepoint detection analysis that are similar to ours, but more limited in scope and usability. Nonparametric methods for changepoint detection are present in the ecp (James, Zhang, and Matteson 2024) and changepoint.np (Haynes and Killick 2022) packages. We consider these beyond our scope at this time, which is focused on parameteric models for changepoint detection. A Bayesian approach is implemented in the bcp (X. Wang, Erdman, and Emerson 2018; Erdman and Emerson 2008) package, which is no longer available on CRAN. The qcc (Scrucca 2017) package is more broadly focused on quality control and statistical process control, neither of which are specific to changepoint detection. The mcp package focuses on regression-based, Bayesian methods. There are many other changepoint packages, all of which have their strengths and limitations. Please see the “Comparison to other packages” vignette in the mcp documentation for a thorough comparison.
@@ 1.5 Our contribution
In this paper, we present several substantive improvements to the existing state-of-the-art. First, the tidychangepoint package (Baumer et al. 2024) provides a unified interface that makes working with existing changepoint detection algorithms in a tidyverse-compliant syntax easy (Section 3). This reduces friction for anyone who wants to compare the results of algorithms, models, or penalized objective functions across different packages. Second, the architecture of the package is easily extended to incorporate new changepoint detection algorithms (Section 4), new parametric models (Section 5), and new penalized objective functions (Section 6). Indeed, our careful separation of these three essential elements of changepoint detection is in and of itself clarifying. Third, to the best of our knowledge, tidychangepoint is the first R package that allows users to run genetic algorithms for changepoint detection without having to write custom code. With GA as a backend, tidychangepoint provides a user-friendly interface for mixing-and-matching various parametric models, penalized objective functions, and initial populations in ways that previously required non-trivial programming by the user. Since tidychangepoint wraps functionality from other packages it adds only modest computational overhead, and delivers correct results (Section 7). Finally, tidychangepoint includes new or updated data sets packaged as time series objects about particulate matter in Bogotá, monthly rainfall in Medellín, temperature in Central England, and batting performance in Major League Baseball. We illustrate the utility of the package in Section 2 and conclude in Section 8.
tidychangepoint is available via download from the Comprehensive R Archive Network (CRAN) since August 19, 2024. The package website contains several short articles demonstrating how to use various features of the package, some of which are not included in this paper (for brevity).
2 Examples using tidychangepoint
In this Section we provide more detailed examples of how tidychangepoint can facilitate meaningful comparisons across algorithms, models, and penalized objective functions.
2.1 Temperatures in Central England
In Section 1.1, we noted that the comparative analysis across different models presented in Shi et al. (2022) required writing custom code for each model fit. In Section 5, we list the many parametric models that tidychangepoint includes. These include the “meanshift” and “trendshift” models, either of which can include AR(1) lagged errors, and either of which can be fit using the BIC or MDL penalties (see Section 6).
With tidychangepoint, we can employ a machine learning style approach to finding the best combination of model and penalty function for a time series. In this case, we build a data.frame with 8 possible combinations.
cet_options <- expand_grid(
  model_fn = c(fit_meanshift_norm, fit_meanshift_norm_ar1,
               fit_trendshift, fit_trendshift_ar1),
  penalty_fn = c(BIC, MDL)
Next, using functionality from the purrr package (Wickham and Henry 2023), we simply iterate the segment() function—which is the main workhorse in tidychangepoint—over these combinations. In each case we employ a genetic algorithm that can produce up to 5000 generations (maxiter), although it will stop if there is no improvement in the fitness after 100 consecutive generations (run).
library(purrr)
cet_mods <- cet_options |>
  mutate(
    tidycpt = map2(
      model_fn, penalty_fn, .f = segment,
      x = CET, method = "ga"maxiter = 5000run = 100
The fitness(), changepoints() and model_name() functions can recover useful information from the resulting objects (see Section 3.2).
cet_mods <- cet_mods |>
  mutate(
    fitness = map_dbl(tidycpt, fitness),
    penalty = map_chr(tidycpt, ~names(fitness(.x))),
    num_cpts = map_int(tidycpt, ~length(changepoints(.x))),
    model_name = map_chr(tidycpt, model_name),,
    num_gens = map_int(tidycpt, ~nrow(as.segmenter(.x)@summary))
We can now easily compare the results and learn about which model performed best with which penalty function.
cet_mods |>
  arrange(penalty, fitness) |>
  select(model_name, fitness, penalty, num_cpts, num_gens)
# A tibble: 8 x 5
  model_name         fitness penalty num_cpts num_gens
  <chr>                <dbl> <chr>      <int>    <int>
1 trendshift            671. BIC            3     1246
2 meanshift_norm_ar1    674. BIC            6     1362
3 meanshift_norm        680. BIC            9     1415
4 trendshift_ar1        682. BIC            4     1347
5 trendshift_ar1        662. MDL            5     2987
6 meanshift_norm        680. MDL            8     1480
7 meanshift_norm_ar1    683. MDL            5     1451
8 trendshift            702. MDL            9     1754
In this case, the trendshift model with white noise errors produces the best fitness score with the BIC penalty, while adding white noise errors produces the best fitness score with the MDL penalty. These results accord with the findings by Shi et al. (2022).
2.2 The designated hitter in Major League Baseball
Sometimes we know the truth about when changepoints occur. For example, Figure 4 shows the difference in batting average between the American and National Leagues of Major League Baseball (MLB) from 1925–2020. The introduction of the designated hitter rule in the American League in 1973 led to noticeable differences in batting performance relative to the National League (which did not adapt the rule until 2020), because the designated hitter employed by American League teams was always a better hitter than the pitcher (who was forced to bat in the National League). Note that in Figure 4, the American League had a higher batting average in every single year from 1973 to 2020. Thus, the year 1973 is a known changepoint, because the data generating process changed before and after that year.
Figure 4: Difference in the rate of home runs per plate appearance between the American and National Leagues in Major League Baseball, 1925–2022. The designated hitter rule was adopted in 1973 by the American League, but not until 2020 by the National League. Note how, during the period from 1973 to 2020 when only the American League employed a designated hitter, the difference in batting average was always negative.
Using the segment() function, we create a list of tidycpt model objects (see Section 3). We apply the PELT algorithm for a change in means and variances using the default MBIC penalty function, and also try a model that allows only the mean to vary across regions and uses the BIC penalty function. Next, we try the WBS algorithm. Finally, we try a genetic algorithm using a normal meanshift model and the MDL penalty function (see Section 6). This algorithm will stop after 500 generations (maxiter), or after 50 consecutive generations with no improvement (run), whichever comes first.
mlb_mods <- list(
  "pelt_meanvar" = segment(mlb_bavg, method = "pelt"),
  "pelt_meanshift" = segment(
    mlb_bavg, method = "pelt",
    model_fn = fit_meanshift_norm, penalty = "BIC"
  "wbs" = segment(mlb_bavg, method = "wbs"),
  "ga" = segment(
    mlb_bavg, method = "ga"model_fn = fit_meanshift_norm,
    penalty_fn = MDL, maxiter = 500run = 50
@@ 2.2 The designated hitter in Major League Baseball
Because each of the resulting objects are of class tidycpt (see Section 3), comparing the results across the four algorithms is easy using map() (from the purrr package (Wickham and Henry 2023)). First, we examine the changepoint sets.
mlb_mods |>
  map(changepoints, use_labels = TRUE|>
  map(as_year)
\$pelt_meanvar
[1] "1976" "2011"
\$pelt_meanshift
character(0)
\$wbs
[1] "1928" "1930" "1942" "1956" "1972"
\$ga
 [1] "1929" "1931" "1936" "1943" "1948" "1952" "1966" "1969" "1973" "1978"
[11] "2012"
PELT finds two changepoints, including 1976, which is just three years after the true changepoint of 1973. It also finds 2011 to be a changepoint, which is interesting because the difference in batting average climbs noticeably around this time. PELT using the meanshift model and the BIC penalty finds no changepoints. WBS finds 5 changepoints, including 1972, just one year before the true changepoint. Our genetic algorithm finds 11 changepoints, including the true changepoint.
To compare other aspects of the fits, we can iterate the glance() function from the broom package (Robinson, Hayes, and Couch 2024) over that same list. Since glance() always returns a one-row tibble, we can then combine those tibbles into one using the list_rbind() function. Care must be taken to ensure that these comparisons are meaningful. While comparing, say, the elapsed time across these algorithms and models is fair, comparing the fitness values is more fraught. MDL scores are not comparable to BIC or MBIC scores, for example.
mlb_mods |>
@@
1 changep~ 2.3     PELT      <list [1]> meanvar    MBIC       -745.  0.011 secs
2 changep~ 2.3     PELT      <list [1]> meanshift~ BIC        -668.  0.002 secs
3 wbs      1.4.1   Wild Bin~ <list [1]> meanshift~ MBIC          0   0.010 secs
4 GA       3.2.4   Genetic   <list [1]> meanshift~ MDL        -735. 19.982 secs
In Figure 5, we compare the changepoint sets found by the four algorithms. While several of the algorithms indicate a changepoint close to the true changepoint of 1973, it is debatable which provides the best fit to the data.
@@
(b) The PELT algorithm using the meanshift model and the BIC penalty.
(c) The WBS algorithm.
(d) A genetic algorithm using the meanshift model and the MDL penalty.
Figure 5: Comparison of changepoint sets returned by four algorithms. Which provides the best fit to the data?
Visually, we could use a rug plot (see ggplot2::geom_rug()) to show the distribution of changepoints across algorithms (see Figure 6). Three of the four algorithms found a changepoint near the true changepoint of 1973. The other changepoints are false positives with respect to the designated hitter rule, but may very well reflect other changes in batting performance that we simply haven’t been thinking about. We suspect that the recent decrease in the difference in batting average could reflect fewer plate appearances taken by pitchers due to the increased use of relief pitchers.
Figure 6: An update to Figure 4 adding a rug for the distribution of changepoints found by the four algorithms mentioned in this section.
2.2.1 Investigating algorithmic performance
Because tidychangepoint provides model-fitting functions and penalized objective functions that “just work”, it is easy to investigate why candidate changepoint sets of interest may not have been chosen by the algorithm.
Because PELT is known to be optimal under mild conditions, all other candidate changepoint sets should return a value for MBIC that is worse than the one found by PELT (which includes the two changepoints 1976 and 2011). The fitness() function gives us this value.
fitness(mlb_mods[["pelt_meanvar"]])
     MBIC
-744.8353
As expected, the MBIC value of the meanvar model with the true changepoint set is worse (further from $\infty$- ∞).
MBIC(fit_meanvar(mlb_bavg, tau = 49))
[1] -700.8384
Why didn’t PELT pick the changepoint set returned by the genetic algorithm? This set, too, had a worse MBIC score, although its score is better than the true changepoint set.
MBIC(fit_meanvar(mlb_bavg, tau = changepoints(mlb_mods[["ga"]])))
[1] -725.6102
Alternatively, we might ask why the genetic algorithm didn’t pick either the true changepoint set or the one found by PELT. Recall that the genetic algorithm is randomized and only a heuristic, and so unlike in the case of PELT, it is possible that we could manually identify a changepoint set with a lower MDL than the one found by the algorithm. However, in this case, the MDL score for the changepoint set returned by the genetic algorithms bests the other two possibilities.
fitness(mlb_mods[["ga"]])
      MDL
-735.2082
list(49changepoints(mlb_mods[["pelt_meanvar"]])) |>
  map(fit_meanshift_norm, x = mlb_bavg) |>
  map_dbl(MDL)
[1] -691.7307 -682.8757
This type of manual investigation is very difficult to do with other changepoint detection packages.
2.3 Particular matter in Bogotá, Colombia
There appears to be some seasonality in the particulate matter data as well as a pronounced drop in particulate matter around the time of the COVID-19 lockdown in March of 2020. Suárez-Sierra, Coen, and Taimal (2023) use Coen’s algorithm (see Section 4.3) with a manually-specified threshold of 37 $\mu g/m^{3}$ informed by Colombian government policy. Here, we reproduce that result using the GA-based implementation in tidychangepoint. Although we find a different changepoint set, the fit of the NHPP model is of comparable quality.
bog_coen <- bogota_pm |>
  segment(
    method = "ga-coen"maxiter = 1000run = 100,
    model_fn_args = list(threshold = 37)
Alternatively, a careful look at the data suggest that the trendshift model, which allows for a linear slope within each region, might be more appropriate. Setting the method argument of segment() to ga allows us to employ a custom genetic algorithm for which we can specify a model via the model_fn argument. Here, we use the MDL() penalty function and the log_gabin_population() function. While Figure 8 picks up on the COVID-19-related changepoint, Figure 8 shows that the trendshift model also fits the data very well.
bog_ga <- bogota_pm |>
  segment(
    method = "ga"model_fn = fit_trendshift, penalty_fn = MDL,
    population = log_gabin_population(bogota_pm),
    maxiter = 500run = 100
(a) Coen’s algorithm.
(b) A genetic algorithm with a trendshift model.
Figure 8: Changepoint sets for two genetic algorithms applied to the Bogotá particulate matter data. Both pick up on the COVID-19-related changepoint in March of 2020. The latter includes linear trends in each region.
fitness(bog_ga)
     MDL
8819.181
bogota_pm |>
  fit_trendshift_ar1(tau = changepoints(bog_ga)) |>
  MDL()
[1] 8791.355
(a) The diagnose() function applied to a model object of class nhpp (Coen’s algorithm).
(b) The diagnose() function applied to a model object of class mod_cpt (a genetic algorithm with a trendshift model).
Figure 9: Diagnostic plot for two algorithms applied to the Bogotá particulate matter data. Note how the diagnostic plots differ based on the model type.